A BODY OF 5 KG RISES BY AN INCLINED PLAIN 30 WITH RESPECT TO THE GROUND , WHEN A FORCE OF 100 IS APPLIED TO IT N THAT PULLS IT CALCULATE THE ACCELERATION WITH WHICH THE BODY ASCENDS TAKING INTO ACCOUNT THAT THERE IS NO FRICTION SUN = 15 , 1 M/S


1
Expert's answer
2021-02-24T16:59:21-0500

"F=ma+mg\sin{30}\\100=5(a+9.8\sin{30})\\a=15.1\frac{m}{s^2}"